# Combating Confirmation Bias: A Unified Pseudo-Labeling Framework for Entity Alignment

### v2

@@ 4.3.2 Effectiveness as a General Pseudo-labeling Framework
4.3.3 Comparisons with Other Pseudo-label Ensembling Methods
[…]
Table 8: Performance of UPL-EA using different pseudo-Label ensembling methods.
DBP15KZH_EN DBP15KJA_EN DBP15KFR_EN
Hit@1 Hit@10 MRR Hit@1 Hit@10 MRR Hit@1 Hit@10 MRR
[…]
To provide deeper insights into the performance of different pseudo-label ensembing methods, we explicitly evaluate the precision of pseudo-labeling using different ensembling methods on the test data, where the ground truth alignments are provided. Fig. 3 compares the precision of pseudo-labeling using three ensembling methods and without ensembling throughout pseudo-labeling iterations.
(a) DBP15KZH_EN
(b) DBP15KJA_EN
(c) DBP15KFR_EN
Figure 3: Comparison with other pseudo-labeling ensembing methods.
Notably, pseudo-labeling with ensembling methods consistently achieves higher precision compared to pseudo-labeling without ensembling throughout the training process. The performance gap is particularly significant on DBP15KZH_EN and DBP15KJA_EN, where larger linguistic barriers exist. Without ensembling, the precision of pseudo-labeling initially increases but quickly plateaus at a sub-optimal level due to confirmation bias arising from the accumulation of one-to-one misalignments (Type II pseudo-labeling errors). In contrast, pseudo-label ensembling maintains a higher precision throughout the training process.
Among ensembling methods, the precision of pseudo-labeling with temporal ensembling initially increases by aggregating the first two iterations but gradually decreases as more iterations are used for ensembling. Majority vote consistently achieves higher precision than temporal ensembling, although its precision is unstable throughout the training process. Overall, consensus ensembling (ours) achieves nearly 100% precision after using only three iterations for ensembling across all KG pairs. Our findings suggest that UPL-EA’s pseudo-label ensembing provides a simple but effective way to improve the quality of pseudo-labeled alignments, achieving competitive performance compared to other ensembling methods.
@@ 4.3.3 Comparisons with Other Pseudo-label Ensembling Methods
UPL-EA 98.3 98.2 98.0 97.8 98.4 98.4 97.7 97.4 -1.0
As expected, UPL-EA consistently outperforms four competitors on all cross-lingual KG pairs at all prior seed alignment rates. This is due to UPL-EA’s ability to augment the training set with reliable pseudo-labeled alignments by effectively alleviating confirmation bias. As the rate of prior seed alignments decreases from 40% to 10%, the performance of BootEA significantly degrades by over 20% on average due to its limited ability to prevent the accumulation of pseudo-labeling errors. RNM outperforms RDGCN owing to its posterior embedding distance editing during pseudo-labeling, however its lack of iterative model re-training hinders its overall performance. CPL-OT demonstrates more stable performance with varying rates of prior seed alignments because it selects pseudo-labeled alignments via the conflict-aware OT modeling and then uses them to train the EA model in turn; nevertheless, its neglect of one-to-one misalignments (Type II pseudo-labeling errors) limits the potential of CPL-OT. UPL-EA remains consistently competitive and stable across all datasets, with an average performance loss of 1% at most when the rate of prior seed alignments decreases from 40% to 10%, even on the most challenging DBP15KZH_EN dataset.
@@ 4.6 Hyper-parameter Sensitivity Analysis
We further study the sensitivity of UPL-EA with regards to four hyper-parameters: embedding dimension $d$, number of iterations $m$ for pseudo-label ensembling, regularization hyper-parameter $\beta$ in Eq. (11), and margin hyper-parameter $\gamma$ in the alignment loss function Eq. (15). This set of sensitivity analysis is conducted on DBP15KZH_EN with 30% prior seed alignments as a case study. The respective results in terms of Hit@1 and Hit@10 are reported in Fig. 4.
@@
(b) Number of iterations for ensembling $m$
@@
As shown in Fig. 4(a), the performance of UPL-EA improves considerably as the embedding dimension $d$ increases from 100 to 300 and then retains at a relatively stable level. For the number of iterations $m$ used for pseudo-label ensembling, the use of ensembling over multiple iterations ($m>1$) significantly improves the alignment performance over a single iteration ($m=1$). This demonstrates the effectiveness of our self-ensembling mechanism, which requires only a few iterations (e.g., $m=3$) to achieve competitive performance (see Fig. 4(b)). In addition, Fig. 4(c) shows that the performance of UPL-EA is insensitive to different values of $\beta$ used in OT-based pseudo-labeling. As for the margin parameter $\gamma$, the performance of UPL-EA begins to drop gradually when $\gamma$ exceeds 1, as shown in Fig. 4(d). This is reasonable, as a larger margin would allow more tolerance for alignment errors, thereby degrading model performance.
@@ 4.7 Runtime Comparison
Lastly, we compare the overall training time of UPL-EA with three representative EA models, CPL-OT, RDGCN and RNM, across five across-lingual datasets with 30% prior seed alignments. For a fair comparison, we use the same parameters reported in the original papers of the three baselines. Fig. 5 reports the overall training time of UPL-EA and the other baselines. Our results show that UPL-EA is considerably more efficient than the three baselines, achieving a speedup of at least 50% across all five datasets.
@@
Notably, CPL-OT and RNM take twice as long as UPL-EA, and three times as long in larger datasets such as DBP15KFR_EN. Additionally, the supervised model RDGCN requires over 60-minute training time on DBP15K and over 30 minutes on SRPRS, indicating its poor runtime efficiency. Overall, our findings suggest that UPL-EA exhibits superior runtime efficiency compared to strong baselines.
@@ 5.1 Entity Alignment in Knowledge Graphs
The entity alignment (EA) task aims to discover one-to-one equivalent entity pairs across two KGs that refer to the same real-world identity. Early EA models are probabilistic methods that compute similarities and perform equivalence reasoning in the input space. For example, PARIS [36] is an unsupervised ontology alignment model that matches entities, relations and entity classes to estimate the likelihood of two entities being equivalent. It is later extended to a semi-supervised EA model, PARIS+ [16], allowing the incorporation of prior seed alignments.
@@
To tackle the shortage of prior seed alignments, semi-supervised EA models have been proposed in recent years. As a prominent learning paradigm among such, pseudo-labeling-based methods, e.g., BootEA [39], IPTransE [61], TransEdge [40], RNM [62], MRAEA [20], and CPL-OT [10], propose to iteratively pseudo-label unaligned entity pairs and add them to prior seed alignments for subsequent model retraining. For RNM, there is a slight difference that it augments prior seed alignments to rectify embedding distance after the completion of model training. Although these methods have achieved promising performance gains, the confirmation bias associated with iterative pseudo-labeling has been largely under-explored. Recent methods like RNM [62] and MRAEA [20] use simple heuristics to preserve only the most convincing alignment pairs, for example, those with the smallest distance, at the presence of conflicts. BootEA [39] and CPL-OT [10], on the other hand, model the inference of pseudo-labeled alignments as an assignment problem, where the most likely aligned pairs are selected at each pseudo-labeling iteration. Unlike BootEA that selects a small set of pseudo-labeled alignments using a pre-specified threshold, CPL-OT imposes a full match between two unaligned entity sets to maximize the number of pseudo-labeled alignments at each iteration. Both methods impose constraints to enforce hard alignments to alleviate alignment conflicts, which may potentially increase one-to-one misalignments.
This work is thus proposed to explicitly address confirmation bias in pseudo-labeling-based entity alignment. We analytically identify two types of pseudo-labeling errors that lead to confirmation bias and propose a new UPL-EA framework to alleviate these errors. Different from our previous work CPL-OT [10], UPL-EA introduces a discrete OT formulation aimed at addressing Type I pseudo-labeling errors. This formulation allows for a more accurate, probabilistic alignment configuration optimized efficiently using the Sinkhorn algorithm. Unlike CPL-OT, which relies on a pre-specified threshold, the threshold for selecting pseudo-labeled alignments in UPL-EA is mathematically derived and proven empirically effective, facilitating its applicability across various datasets. In addition, a self-ensembling approach is further proposed to refine pseudo-labeled alignments by combining predictions over multiple pseudo-labeling iterations, thus mitigating Type II pseudo-labeling errors.
@@ 5.2 Pseudo-Labeling
Pseudo-labeling has emerged as an effective semi-supervised approach in addressing the challenge of label scarcity. It refers to a self-training paradigm where the model is iteratively bootstrapped with additional labeled data based on its own predictions. The pseudo-labels generated from model predictions can be defined as hard (one-hot distribution) or soft (continuous distribution) labels [15, 33, 1]. More specifically, pseudo-labeling strategies are designed to select high-confidence unlabeled data by either directly taking the model’s predictions, or sharpening the predicted probability distribution. It is closely related to entropy regularization [32], where the model’s predictions are encouraged to have low entropy (i.e., high-confidence) on unlabeled data. The selected pseudo-labeleds are then used to augment the training set and to fine-tune the model initially trained on the given labels. This training regime is also extended to an explicit teacher-student configuration [29], where a teacher network generates pseudo-labeleds from unlabeled data, which are used to train a student network.
@@ 5.3 Optimal Transport on Graphs
Optimal Transport (OT) is the general problem of finding an optimal plan to move one distribution of mass to another with the minimal cost [47]. As an effective metric to define the distance between probability distributions, OT has been applied in computer vision and natural language processing over a range of tasks including machine translation, text summarization, and image captioning [46, 6]. In recent years, OT has also been studied on graphs to match graphs with similar structures or align nodes/entities across graphs. For graph partitioning and matching, the transport on the edges across graphs is used to define the Gromov-Wasserstein (GW) discrepancy [45] that measures how edges in a graph compare to those in another graph [55, 22, 54]. For entity alignment across graphs, Pei et al [27] incorporate an OT objective into the overall loss to enhance the learning of entity embeddings. Tang et al [42] propose to jointly perform structure learning and OT alignment through minimizing multi-view GW distance matrices between two attributed graphs. These methods have primarily used OT to define a learning objective, which involves bi-level optimization for model training. To further enhance the scalability of OT modeling for entity alignment, Mao et al [21] propose to make the similarity matrix sparse by dropping its entries close to zero. However, this sparse OT modeling potentially violates the constraints of the OT objective, failing to guarantee one-to-one correspondences across two KGs. In our work, we focus on tackling the scarcity of prior seed alignments via iterative pseudo-labeling; we seek to find more accurate one-to-one alignment configurations between entities via OT modeling. thus eliminating conflicted misalignments at each pseudo-labeling iteration and mitigating confirmation bias.
@@ 6 Conclusion and Future Work
We have investigated the problem of confirmation bias for pseudo-labeling-based entity alignment, which has been largely overlooked in the literature. Through an in-depth analysis, we have revealed the underlying causes of confirmation bias and proposed a novel unified pseudo-labeling framework (UPL-EA) for entity alignment. UPL-EA systematically addresses confirmation bias through two key innovations: OT-based pseudo-labeling and pseudo-label ensembling. OT-based pseudo-labeling utilizes a discrete OT formulation to more accurately infer pseudo-labled alignments that satisfy one-to-one correspondences, thus mitigating Type I pseudo-labeling errors within each iteration. Pseudo-label ensembling uses a self-ensembling approach that combines the predictions of pseudo-labeled alignments over multiple pseudo-labeling iterations to reduce pseudo-label selection variability, thus alleviating the propagation of Type II pseudo-labeling errors into subsequent model training. Our extensive experimental evaluation and analysis demonstrate that UPL-EA outperforms state-of-the-art baselines across various types of benchmark datasets. The competitive performance of UPL-EA validates its superiority in addressing confirmation bias and its utility as a general pseudo-labeling framework to improve entity alignment performance. Future research will include a theoretical investigation to rigorously assess the effectiveness of pseudo-labeling ensembling within UPL-EA. Additionally, we will explore extending our OT formulation to incorporate different sources of information for multi-modal entity alignment.

### v3

@@ 4.3.3 Effectiveness as a General Pseudo-Labeling Framework
UPL-EA 98.2 98.2 98.0 97.9 98.4 98.4 97.7 97.4 -0.7
As expected, UPL-EA consistently outperforms four competitors on all cross-lingual KG pairs at all prior seed alignment rates. This is due to UPL-EA’s ability to augment the training set with reliable pseudo-labeled alignments by effectively alleviating confirmation bias. As the rate of prior seed alignments decreases from 40% to 10%, the performance of BootEA significantly degrades by over 20% on average due to its limited ability to prevent the accumulation of pseudo-labeling errors. RNM outperforms RDGCN owing to its posterior embedding distance editing during pseudo-labeling; however, its lack of iterative model re-training hinders its overall performance. CPL-OT demonstrates more stable performance with varying rates of prior seed alignments because it selects pseudo-labeled alignments via the conflict-aware OT modeling and then uses them to train the EA model in turn; nevertheless, its neglect of one-to-one misalignments limits the potential of CPL-OT. UPL-EA remains consistently competitive and stable across all datasets, with an average performance loss of 1% at most when the rate of prior seed alignments decreases from 40% to 10%, even on the most challenging DBP15KZH_EN dataset.
@@ 4.6 Hyper-Parameter Sensitivity Analysis
We further study the sensitivity of UPL-EA with regards to four hyper-parameters: embedding dimension $d$, number of OT-based models $M$ for pseudo-label ensembling, regularization hyper-parameter $\beta$ in Eq. (8), and margin hyper-parameter $\gamma$ in the alignment loss function Eq. (2). This set of sensitivity analysis is conducted on DBP15KZH_EN with 30% prior seed alignments as a case study. The respective results in terms of Hit@1 and Hit@10 are reported in Fig. 3.
@@
(b) Number of models for ensembling $M$
@@
As shown in Fig. 3(a), the performance of UPL-EA improves considerably as the embedding dimension $d$ increases from 100 to 300 and then retains a relatively stable level. For the number of OT-based models $M$ used for parallel pseudo-label ensembling, the use of ensembling over multiple OT-based models ($M>1$) significantly improves the alignment performance over a single one ($M=1$). This demonstrates the effectiveness of our parallel ensembling mechanism, which requires only a few OT-based models (e.g., $M=3$) to achieve competitive performance (see Fig. 3(b)). In addition, Fig. 3(c) shows that the performance of UPL-EA is insensitive to different values of $\beta$ used in OT-based pseudo-labeling. As for the margin parameter $\gamma$, the performance of UPL-EA begins to drop gradually when $\gamma$ exceeds 1, as shown in Fig. 3(d). This is reasonable, as a larger margin would allow more tolerance for alignment errors, thereby degrading model performance.
@@ 4.7 Runtime Comparison
Lastly, we compare the overall training time of UPL-EA with three embedding-based EA models, including CPL-OT, RDGCN, and RNM, and one conventional EA model, PARIS+, across five cross-lingual datasets with 30% prior seed alignments. For a fair comparison, we use the same parameters reported in the original papers of the four baselines. Fig. 4 reports the overall training time of UPL-EA and the other baselines. Our results show that UPL-EA is considerably more efficient than the three embedding-based baselines, achieving a speedup of at least 50% across all five datasets. Although PARIS+ exhibits the shortest runtime due to its rule-based nature and lack of gradient-based optimization, UPL-EA remains highly efficient while maintaining strong EA performance.
Notably, CPL-OT and RNM take more than twice as long as UPL-EA, and three times as long on larger datasets such as DBP15KFR_EN. Additionally, the supervised model RDGCN requires over 60-minute training time on DBP15K and almost 30 minutes on SRPRS, indicating its poor runtime efficiency. Overall, our findings suggest that UPL-EA exhibits superior runtime efficiency compared to strong embedding-based baselines, while maintaining a well-balanced trade-off between EA performance and time efficiency compared to conventional EA methods.
@@ 5.1 Entity Alignment in Knowledge Graphs
The entity alignment (EA) task aims to discover one-to-one equivalent entity pairs across two KGs that refer to the same real-world identity. Early EA models are probabilistic methods that compute similarities and perform equivalence reasoning in the input space. For example, PARIS [37] is an unsupervised ontology alignment model that jointly aligns entities, relations, and classes across KGs. It converts each relation into a logic-rule based function, and iteratively refines entity alignment probabilities using the functional and inverse functional properties of relations. PARIS is later extended to a semi-supervised EA model, PARIS+ [17], allowing the incorporation of prior seed alignments.
@@
To tackle the shortage of prior seed alignments, semi-supervised EA models have been proposed in recent years. As a prominent learning paradigm among such, pseudo-labeling-based methods, e.g., BootEA [40], IPTransE [62], TransEdge [41], RNM [63], MRAEA [21], and CPL-OT [10], propose to iteratively pseudo-label unaligned entity pairs and add them to seed alignments for subsequent model training. For RNM, there is a slight difference that it augments seed alignments to rectify embedding distance after the completion of model training. Although these methods have achieved promising performance gains, the confirmation bias associated with iterative pseudo-labeling has been largely under-explored. Recent methods like RNM [63] and MRAEA [21] use simple heuristics to preserve only the most convincing alignment pairs, for example, those with the smallest distance, at the presence of conflicts. BootEA [40] and CPL-OT [10], on the other hand, model the inference of pseudo-labeled alignments as an assignment problem, where the most likely aligned pairs are selected at each pseudo-labeling iteration. Unlike BootEA that selects a small set of pseudo-labeled alignments using a pre-specified threshold, CPL-OT imposes a full match between two unaligned entity sets to maximize the number of pseudo-labeled alignments at each iteration. Both methods impose constraints to enforce hard alignments to alleviate alignment conflicts, which may potentially increase one-to-one misalignments.
This work is thus proposed to explicitly address confirmation bias in pseudo-labeling-based entity alignment. We analytically identify two types of pseudo-labeling errors that lead to confirmation bias and propose a new UPL-EA framework to alleviate these errors. Different from our previous work CPL-OT [10], UPL-EA introduces a discrete OT formulation aimed at addressing conflicted misalignments. This formulation allows for a more accurate, probabilistic alignment configuration optimized efficiently using the Sinkhorn algorithm. Unlike CPL-OT, which relies on a pre-specified threshold, the threshold for selecting pseudo-labeled alignments in UPL-EA is mathematically derived and proven empirically effective, facilitating its applicability across various datasets. In addition, a parallel ensembling approach is further proposed to refine pseudo-labeled alignments by combining predictions over multiple OT-based models trained in parallel, thus mitigating one-to-one misalignments.
@@ 5.2 Pseudo-Labeling
Pseudo-labeling has emerged as an effective semi-supervised approach in addressing the challenge of label scarcity. It refers to a self-training paradigm where the model is iteratively bootstrapped with additional labeled data based on its own predictions. The pseudo-labels generated from model predictions can be defined as hard (one-hot distribution) or soft (continuous distribution) labels [16, 34, 1]. More specifically, pseudo-labeling strategies are designed to select high-confidence unlabeled data by either directly taking the model’s predictions, or sharpening the predicted probability distribution. It is closely related to entropy regularization [33], where the model’s predictions are encouraged to have low entropy (i.e., high-confidence) on unlabeled data. The selected pseudo-labels are then used to augment the training set and to fine-tune the model initially trained on the given labels. This training regime is also extended to an explicit teacher-student configuration [30], where a teacher network generates pseudo-labels from unlabeled data, which are used to train a student network.
@@ 5.3 Optimal Transport on Graphs
Optimal Transport (OT) is the general problem of finding an optimal plan to move one distribution of mass to another with the minimal cost [48]. As an effective metric to define the distance between probability distributions, OT has been applied in computer vision and natural language processing over a range of tasks including machine translation, text summarization, and image captioning [47, 6]. In recent years, OT has also been studied on graphs to match graphs with similar structures or align nodes/entities across graphs. For graph partitioning and matching, the transport on the edges across graphs is used to define the Gromov-Wasserstein (GW) discrepancy [46] that measures how edges in a graph compare to those in another graph [56, 23, 55]. For entity alignment across graphs, Pei et al [28] incorporate an OT objective into the overall loss to enhance the learning of entity embeddings. Tang et al [43] propose to jointly perform structure learning and OT alignment through minimizing multi-view GW distance matrices between two attributed graphs. These methods have primarily used OT to define a learning objective, which involves bi-level optimization for model training. To further enhance the scalability of OT modeling for entity alignment, Mao et al [22] propose to make the similarity matrix sparse by dropping its entries close to zero. However, this sparse OT modeling potentially violates the constraints of the OT objective, failing to guarantee one-to-one correspondences across two KGs. In our work, we focus on tackling the scarcity of prior seed alignments via iterative pseudo-labeling; we seek to find more accurate one-to-one alignment configurations between entities via OT modeling, thus eliminating conflicted misalignments at each pseudo-labeling iteration and mitigating confirmation bias.
@@ 6 Conclusion and Future Work
We have investigated the problem of confirmation bias for pseudo-labeling-based entity alignment, which has been largely overlooked in the literature. Through an in-depth analysis, we have revealed the underlying causes of confirmation bias and proposed UPL-EA, a novel unified pseudo-labeling framework for entity alignment. UPL-EA systematically addresses confirmation bias through two key innovations: OT-based pseudo-labeling and parallel pseudo-label ensembling. OT-based pseudo-labeling utilizes a discrete OT formulation to more accurately infer pseudo-labeled alignments that satisfy one-to-one correspondences, thus mitigating conflicted misalignments. Parallel pseudo-label ensembling combines the predictions of pseudo-labeled alignments from multiple OT-based models independently trained in parallel to reduce variability in pseudo-label selection, thus alleviating the propagation of one-to-one misalignments into subsequent model training. Our extensive experimental evaluation and analysis demonstrate that UPL-EA outperforms state-of-the-art baselines across various types of benchmark datasets. The competitive performance of UPL-EA validates its superiority in addressing confirmation bias and its utility as a general pseudo-labeling framework to improve entity alignment performance. Future research will include a theoretical investigation to rigorously assess the effectiveness of pseudo-labeling ensembling within UPL-EA. Additionally, we will explore extending our OT formulation to incorporate different sources of information for multi-modal entity alignment.